OpenStudy (anonymous):

factor polynomial x^2-x-42

OpenStudy (anonymous):

Find two numbers that add up to -1 but multiply to equal -42.

OpenStudy (anonymous):

x-42 is the factored equation

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

x^2 -x -42 You have to find two numbers that add up to be -1 and multiply to equal -42 -7 * 6 = -42 -7 + 6 = -1 Thus: (x + 6)(x - 7) is the factored form. Feel free to multiple (x + 6)(x - 7) to see.

OpenStudy (anonymous):

@Speedy_Jazz